OpenStudy (liizzyliizz):

h(x)= sin2xcos2x i know the rules to find the derivative but idk how the answer is 2cos4x. help?

OpenStudy (amistre64):

so what do you end at?

OpenStudy (amistre64):

is that sin(2x) cos(2x) or are those squared?

OpenStudy (liizzyliizz):

no just 2.

OpenStudy (amistre64):

then product and pop out the chains

OpenStudy (amistre64):

sin(2x)' cos(2x) +sin(2x) cos(2x)' 2 cos(2x) cos(2x) -2 sin(2x)sin(2x) 2 (cos(2x) cos(2x) - sin(2x)sin(2x)) what do you remember from trig identities?

OpenStudy (liizzyliizz):

i know u for the chain rule means expressions, but my teacher didnt explain it at all today so im confused with it. what do u with it o.o

OpenStudy (amistre64):

do you mean, what is the chain rule? its like peeling an onion so that all the parts are accounted for

OpenStudy (amistre64):

the trig identity they use is: cos (a+b) = cos(a) cos(b) - sin(a) sin(b) cos(2x + 2x) = cos(4x) right? so if anything its the chain rule that you might have an issue with

OpenStudy (liizzyliizz):

im sorry im going thru my old notes from trig identities from last year lol. AND OOOH. god i wish i had memorized these identities better =/

OpenStudy (liizzyliizz):

well i just learned about it today, my teacher just showed us it, didnt tell us what to do with it, and im here doing hw not knowing how to apply it. so im kind of learning this right now lol